UT20090705 LUCIFER Commissioning

Observers: WSeifert, NAgeorges (LBT), DThompson (Tucson)
Telescope Operator: D Gonzalez Huerta
Software Support: CBiddick (Tucson)
Instrument Support:
Telescope Support:

Summary

  • Tucson TO station having problems...notes added to IT#1575, priority=high

Details (times in UT unless otherwise noted)

Opened dome ~03:00 (20:00 MST)! Clouds thinning. Seeing not great at ~1.5 arcsec FWHM on guider.

GCS issuing warnings more frequently than necessary. We still need to work on the thresholding, exposure time scaling, and warning messages.

~02:00 The guideCorrectionFactor was found to be 0.9 in AIP_L.cfg, changed to 0.75. Last month we were using AIP_IRTC_L.cfg.

02:41 Toutside=11.8C Tmirror=14.6C RH=62.5% wind=6.5 peak

05:38 With the LUCIFER calibration unit in the beam, GCS mostly correctly failed to find a guide star and thus did not send offsets to PCS. But occasionally it was apparently detecting something:
Jul  5 05:39:23 lbtmu304 LBT_left_GCS: no star found. peak: 1287.000000, mbg: 1258.963320
Jul  5 05:39:25 lbtmu304 LBT_left_GCS: no star found. peak: 1287.000000, mbg: 1259.457441
Jul  5 05:39:28 lbtmu304 LBT_left_GCS: no star found. peak: 1287.000000, mbg: 1259.039188
Jul  5 05:39:31 lbtmu304 LBT_left_GCS: centx: 233.000000, centy: 327.000000, xf: 0.000000, yf: 97.000000,peak: 1281.000000, mbg: 1259.522360
Jul  5 05:39:32 lbtmu304 LBT_left_GCS: PROFILING: sent guide correction to PCS.
Jul  5 05:39:34 lbtmu304 LBT_left_GCS: no star found. peak: 1287.000000, mbg: 1258.781486
Jul  5 05:39:37 lbtmu304 LBT_left_GCS: no star found. peak: 1287.000000, mbg: 1258.701337
Jul  5 05:39:40 lbtmu304 LBT_left_GCS: no star found. peak: 1287.000000, mbg: 1259.094472
If these are really sending offsets to PCS, then we need to stop this from happening. The peak flux numbers look like they should fall below the 5.0*stddev threshold. IT#2261 submitted

Chris found ~1000 defunct processes. He traced them to the PMCGUIs (each had 507 defunct processes attached to them). Killing the PCSGUI with "kill -9" made the defunct processes go away, and the GUIs were restarted without issues.

There is also a java_vm running that is taking up 100% of a CPU (may be multiple threads split across multiple CPUs). We asked DGH to stop his firefox window in which the radar loop was running. This made the java_vm process disappear, so this is the source of the high CPU use.

06:24 DGH reports RH=90% and climbing.

06:48 Closing dome. RH=94% and still rising.

07:50 Re-opened.
07:58 Re-closed. Rising humidity

-- DavidThompson - 05 Jul 2009
I Attachment Action Size Date Who Comment
daq.20090705.gifgif daq.20090705.gif manage 42 K 08 Jul 2009 - 17:51 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r4 - 08 Jul 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback